OPERATIONS SYNOPSIS (continued)Transit to Site 1171
An APC/XCB BHA was run to near the PDR depth of 2176.4 mbrf and Hole 1171A was spudded with the APC at 1415 hr on 12 April. The seafloor depth was calculated to be 2159.4 mbrf, or 2148.2 mbsl, based on recovery of the first core. Piston coring advanced to 111.6 mbsf (mean recovery = 98%) where Core 12H could not be retrieved from the sediment with 80 Kips overpull, requiring drilling over the core barrel to release it from the sediments. Piston cores were oriented starting with Core 3H and continued through to Core 12H. Heat-flow measurements were obtained with the Adara cutting shoe at 35.6 mbsf (Core 5H) and 54.6 mbsf (Core 7H).
The XCB system was used to deepen the hole to 124.4 mbsf, with 62% recovery. The original operational plan called for the hole to be cored as deeply as possible with the XCB. Because of unusually calm sea conditions, we decided to stop XCB coring after two cores and to continue piston coring the remaining two of the three planned APC holes to potentially avoid weather dependent core disturbance that we had previously encountered.
The vessel was offset 20 m north of Hole 1171A. To obtain a stratigraphic overlap with the cores recovered from Hole 1171A, the bit was positioned 3 m shallower, and Hole 1171B was spudded with the APC at 0530 hr on 13 April. The estimated seafloor depth of 2159.2 mbrf (2148 mbsl) was calculated based on recovery of the first core. Piston coring advanced to 108.8 mbsf with an average recovery of 98%, and cores were oriented starting with Core 3H. The bit was pulled clear of the seafloor at 1545 hr on 13 April, ending Hole 1171B.
The vessel was offset 20 m to the east, and Hole 1171C was spudded with the APC at 1630 hr. To maintain appropriate stratigraphic overlap with the two previous holes, the bit was lowered 6 m from the spud-in depth of Hole 1171B. The seafloor depth was estimated to be 2159.0 mbrf, or 2147.8 mbsl. Piston coring advanced without incident to the depth objective of 104.5 mbsf, with an average recovery of 101%. Cores were oriented beginning with Core 3H and one Adara deployment was made at 38.0 mbsf (Core 5H).
The XCB system was used to deepen Hole 1171C from 104.5 to 274.8 mbsf, where a glauconitic sandstone was encountered. When the ROP slowed to 1.5 m/hr, the refusal depth of the XCB system was declared and coring terminated. The average recovery of the XCB portion of this hole was 82%, with a total average recovery of 89% for Hole 1171C.
Following the recovery of the last core from Hole 1171C, the drill string was recovered, a RCB BHA was made up, and the ship was offset 20 m east of Hole 1171C. Hole 1171D was spudded at 0400 hr on 15 April and drilled ahead with a center bit to 100 mbsf. The center bit was recovered by wireline and the Davis-Villinger Temperature Probe (DVTP) deployed to obtain a temperature measurement at this depth. Following the recovery of the DVTP, the center bit was dropped again and drilling resumed to the depth objective of 247.6 mbsf. The average ROP for the drilled interval was 62 m/hr, excluding the time expended on the DVTP measurement. At 1235 hr on 15 April, RCB coring was initiated and advanced with increasing recovery to the depth objective of 958.8 mbsf. The average recovery for this hole was 74%, with an average ROP of 17 m/hr for the 711-m cored interval. There were no stability problems encountered while drilling and coring in this hole.
Following an 11-hr wiper trip and hole preparation that included the release of the bit at the
bottom of the hole, the pipe was raised to the logging depth of 162.2 mbsf. Three tool string runs
were planned for this hole; the triple combo, the GHMT-sonic, and the FMS. The first logging run
with the triple combo reached the target depth of 958.8 mbsf, and excellent quality data were
collected from the entire section of open hole. While running in with the second log, the GHMT
sonic, the tool string bridged out at ~291 mbsf, near the transition from nannofossil chalk to
Eocene glauconitic sandstones. Apparently while running in with this logging suite, power was
lost to the GHMT-sonic tool when the unit became caught on this bridge. The tool string was
returned to the rig floor, where ~600 m of damaged logging cable was removed and the logging
line was re-terminated.
In an attempt to clear the bridge in the formation, the drill string was lowered with the objective
of clearing all hole obstructions to ~300 mbsf. Perhaps as a result of the 4- to 5.5-m heave, the
open end of the BHA became plugged with sediment after traversing <1 stand of pipe at 190
mbsf. The drill string was worked free with 150,000 lb of additional tension after working the drill
string for 45 min. Because the bottom of the pipe was plugged and we thought that the drill string
had apparently meandered into the side of the borehole, no further logging could be undertaken.
While we retrieved the drill string, the primary beacon was recovered. The alternate beacon was also successfully released; however, the flasher apparently did not activate upon surfacing and the beacon drifted off into the dark and could not be located. The drilling equipment was secured by 0645 hr on 22 April as the ship departed for the next site. The errant beacon was observed on the surface as the vessel left location, but when the ship came about to recover the unit, it could not be found in the 4.5-m swell.
Operations Synopsis-Site 1172 | Table of Contents